A consulting agency was asked by a large insurance company to investigate if business majors were better salespersons than those with other majors. A sample of 20 salespersons with a business degree showed that they sold an average of 11 insurance policies per week. Another sample of 25 salespersons with a degree other than business showed that they sold an average of 9 insurance policies per week. Assume that the two populations are approximately normally distributed with population standard deviations of 1.80 and 1.35 policies per week, respectively. Construct a 99% confidence interval for the difference between the two population means.
Answer #1

99% confidence interval is

From Z table, Z/2 at 0.01 significance level = 2.576

Lower limit = (11-9) - 2.576 * sqrt [ 1.82 / 20 + 1.352 / 25 ] = 0.7515

Upper limit = (11-9) + 2.576 * sqrt [ 1.82 / 20 + 1.352 / 25 ] = 3.2485

99% CI Is ( 0.7515 , 3.2485 )